### v1

# Decoupled Global-Local Alignment for Improving Compositional Understanding

Abstract.
Contrastive Language-Image Pre-training (CLIP) has achieved success on multiple downstream tasks by aligning image and text modalities. However, the nature of global contrastive learning limits CLIP’s ability to comprehend compositional concepts, such as relations and attributes. Although recent studies employ global hard negative samples to improve compositional understanding, these methods significantly compromise the model’s inherent general capabilities by forcibly distancing textual negative samples from images in the embedding space. To overcome this limitation, we introduce a Decoupled Global-Local Alignment (DeGLA) framework that improves compositional understanding while substantially mitigating losses in general capabilities. To optimize the retention of the model’s inherent capabilities, we incorporate a self-distillation mechanism within the global alignment process, aligning the learnable image-text encoder with a frozen teacher model derived from an exponential moving average. Under the constraint of self-distillation, it effectively mitigates the catastrophic forgetting of pretrained knowledge during fine-tuning. To improve compositional understanding, we first leverage the in-context learning capability of Large Language Models (LLMs) to construct about 2M high-quality negative captions across five types. Subsequently, we propose the Image-Grounded Contrast (IGC) loss and Text-Grounded Contrast (TGC) loss to enhance vision-language compositionally. Extensive experimental results demonstrate the effectiveness of the DeGLA framework. Compared to previous state-of-the-art methods, DeGLA achieves an average enhancement of 3.5% across the VALSE, SugarCrepe, and ARO benchmarks. Concurrently, it obtains an average performance improvement of 13.0% on zero-shot classification tasks across eleven datasets. Our code will be released at https://github.com/xiaoxing2001/DeGLA.
Vision-Language Model, Multi-Modal, Compositional Understanding
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; April; 2025isbn: 978-1-4503-XXXX-X/YY/MMccs: Computing methodologies Computer vision representations
(a) General performance.
(b) Compositional performance
(c) An illustration for the compositional understanding.
Figure 1. (a) General performance comparison across 11 classification datasets. (b) Compositional performance comparison on VALSE, ARO, and SugarCrepe benchmarks. (c) CLIP scores for images with aligned and unaligned captions. DeGLA demonstrates significantly enhanced capabilities in compositional understanding.
1. Introduction
The rapid expansion of mobile networks and social platforms significantly boosts the large-scale generation of image-text pairs, laying a crucial foundation for vision-language pre-training. In particular, Contrastive Language-Image Pre-training (Radford et al., 2021a) (CLIP) leverages two distinct unimodal encoders for images and texts, and employs contrastive loss (Oord et al., 2018a) for representation learning. After pretraining on extensive image-text pairs sourced from the internet, CLIP exhibits strong transferability and is extensively applied across various tasks, including image captioning (Alayrac et al., 2022; Li et al., 2023b; Mañas et al., 2022), object detection (Kaul et al., 2023; Ma et al., 2023; Wu et al., 2024b), and segmentation (Li et al., 2022b; Cho et al., 2024; Yu et al., 2023; Minderer et al., 2022; Hoyer et al., 2025; Gu et al., 2021; Zang et al., 2022; Wu et al., 2024a).
As CLIP attracts increasing attention from researchers, several enhanced methods based on CLIP have been proposed (Li et al., 2023a; Gu et al., 2024; Zhang et al., 2024b). SLIP (Mu et al., 2022) introduces a multitask learning framework that integrates self-supervised learning with CLIP pretraining. FILIP (Yao et al., 2021) enhances the fine-grained alignment between image patches and textual words by refining the contrastive loss, while maintaining the capability for offline pre-computation of image and text representations during inference, ensuring efficiency in both large-scale training and inference phases. ALIP (Yang et al., 2023) proposes the use of synthetic captions and adaptive contrastive loss to mitigate the influence of noisy data and improve the efficacy of pre-training data utilization. SigLIP (Zhai et al., 2023) utilizes a sigmoid loss, which supports scaling up batch sizes and also performs effectively at smaller batch sizes. However, global contrastive learning inadequately leverages compositional structures within image-text pairs, thereby limiting CLIP’s capability to capture nuanced compositional information in multimodal data (Yuksekgonul et al., 2022). As shown in Figure 1(c), CLIP struggles to discern the relationship between “athlete” and “basketball”.
Recent studies (Yuksekgonul et al., 2022; Huang et al., 2024a; Zhang et al., 2024a; Doveh et al., 2023a, b) have sought to improve the compositional understanding of CLIP. NegCLIP (Yuksekgonul et al., 2022) introduces the Attribution, Relation, and Order (ARO) benchmark and, for the first time, proposes a fine-tuning framework that integrates hard negative samples to enhance CLIP’s compositional understanding. Structure-CLIP (Huang et al., 2024a) integrates Scene Graph Knowledge (SGK) to augment multimodal structured representations. Hard-positives (Kamath et al., 2024) incorporates hard positive samples during fine-tuning to strengthen the model’s capacity to capture subtle but semantically related variations among similar instances. CE-CLIP (Zhang et al., 2024a) introduces a simple yet effective fine-tuning framework with two fine-grained alignment losses to enhance compositional understanding, achieving state-of-the-art performance on multiple vision-language compositional reasoning benchmarks. However, our analysis reveals that existing methods improve CLIP’s compositional understanding at the cost of general capabilities, exhibiting significant catastrophic forgetting of pre-trained knowledge.
To address the limitation of previous works, we innovatively introduce a Decoupled Global-Local Alignment (DeGLA) framework, which aims to improve compositional understanding while substantially mitigating losses in general capabilities. To optimize the retention of the inherent capabilities of the model, we incorporate a self-distillation mechanism within the global alignment process, aligning the learnable image-text encoder with a teacher model derived from an exponential moving average. To improve compositional understanding, we first leverage the in-context learning capability of Large Language Models (LLMs) to construct about 2M high-quality negative captions across five types. Subsequently, we propose the Image-Grounded Contrast (ICC) and Text-Grounded Contrast (TCC) to improve fine-grained understanding. We conduct extensive experiments and demonstrate that our method achieves new state-of-the-art performance in both compositional and general tasks. The main contributions are summarized as follows:
  • We observe that previous methods, while enhancing CLIP’s compositional understanding, often compromise its general understanding capabilities.
  • We propose a simple yet effective negative caption generation pipeline that leverages the context learning capability of Large Language Models (LLMs) to generate high-quality negative captions.
  • We introduce the DeGLA framework, which employs a self-distillation mechanism within the global alignment to maintain the model’s inherent general comprehension capabilities. Additionally, it combines Image-Grounded Contrast (IGC) loss and Text-Grounded Contrast (TGC) loss to improve vision-language compositional understanding.
  • We conduct extensive experiments and demonstrate that DeGLA achieves new state-of-the-art performance in both compositional and general tasks.
2. Related work
2.1. Vision-Language Contrastive Learning
Vision-language representation learning aims to develop robust representations by pretraining models on extensive image-text pair datasets. A key development in this field, CLIP (Radford et al., 2021b), pretrained on 400 million web-collected image-text pairs using contrastive learning, demonstrates strong zero-shot generalization across a wide range of vision tasks. Recent advancements include several methodologies that build on CLIP’s framework. UniCLIP (Lee et al., 2022) improves data efficiency by integrating contrastive losses from various domains into a single universal space. HiCLIP (Geng et al., 2023) incorporates hierarchy-aware attention mechanisms in both visual and linguistic branches of CLIP, significantly enhancing cross-modal alignment. LaCLIP (Fan et al., 2023) employs large language models to rewrite text, thereby increasing sentence structure and vocabulary diversity while preserving essential concepts and meanings. Nonetheless, these developments do not rectify the limitations in CLIP’s compositional understanding, constrained by the inherent shortcomings of global contrastive learning (Yuksekgonul et al., 2022).
2.2. Vision-Language Compositionality
Recently, there have been some works (Yuksekgonul et al., 2022; Hsieh et al., 2024; Parcalabescu et al., 2021; Thrush et al., 2022; Kamath et al., 2023; Peng et al., 2024) aim to enhance the compositional understanding of CLIP. NegCLIP (Yuksekgonul et al., 2022) introduces the Attribution, Relation, and Order (ARO) benchmark and utilizes hard negatives, consisting of nearest neighboring images within each batch, to force models to discern fine-grained differences in highly similar scenes. Structure-CLIP (Huang et al., 2024a) incorporates Scene Graph Knowledge (SGK) to enhance multimodal structured representations. CE-CLIP (Zhang et al., 2024a) proposes a simple yet effective strategy to optimize the utilization of existing image-text datasets, achieving state-of-the-art performance across multiple vision-language compositional reasoning benchmarks. However, although these methods significantly enhance CLIP’s compositional understanding, they frequently compromise its original generalization capabilities. Developing an approach that simultaneously improves both general and compositional understanding continues to be a substantial challenge.
2.3. Knowledge Distillation
Knowledge distillation (Hinton, 2015) is extensively utilized in various domains, including vision (Huang et al., 2024b; Li et al., 2021, 2023c) and natural language processing (Jiao et al., 2019). Recently, a series of knowledge distillation methods tailored specifically for CLIP have been proposed. TinyCLIP (Wu et al., 2023) introduces affinity mimicking, which explores the interaction between modalities during distillation, enabling student models to replicate the teacher’s behavior in learning cross-modal feature alignment within a visual-linguistic affinity space. CLIP-KD (Yang et al., 2024) proposes several distillation strategies, including relation, feature, gradient, and contrastive paradigms, aimed at maximizing feature similarity between the teacher and student models. CLIP-CID (Yang et al., 2024) employs cluster-instance discrimination to facilitate knowledge transfer from the teacher model to the student model, thereby enabling the student model to develop a comprehensive semantic understanding of the pre-training data. Different from the above method, this paper introduces a self-distillation mechanism in global alignment to mitigate catastrophic forgetting during training and preserve the generalization capabilities of the model.
3. Method
Figure 2. The overview of our proposed LLM-driven negative caption generation pipeline. We leverage the robust in-context learning capabilities of LLM to generate five types of hard negative captions.
In subsequent sections, we first present the preliminary knowledge of CLIP in Section 3.1, followed by a detailed exposition of our proposed LLM-driven negative caption generation pipeline in Section 3.2. A comprehensive description of the DeGLA training framework is provided in Section 3.3.
3.1. Preliminaries of CLIP
CLIP consists of a text encoder $\mathcal{E}_{T}$, and an image encoder $\mathcal{E}_{I}$. It encodes a batch of image-text pairs $\{(I_{i},T_{i})\}_{i=1}^{\mathcal{B}}$ into the feature space $\{(v_{i},t_{i})\}_{i=1}^{\mathcal{B}}$. The model employs the InfoNCE loss (Oord et al., 2018b) to optimize the embeddings by minimizing the distance between corresponding image and text features while maximizing the distance between non-corresponding pairs. The image-to-text contrastive loss function is defined as follows:
(1) $\mathcal{L}_{I\rightarrow T}=-\log\frac{\exp(v_{i}\cdot t_{i}^{\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(v_{i}\cdot t_{j}^{\top}/\tau)},$
where $v_{i}$ serves as the anchor. Given $t_{i}$ as the anchor, The symmetric text-to-image contrastive loss is formulated as follows:
(2) $\mathcal{L}_{T\rightarrow I}=-\log\frac{\exp(t_{i}\cdot v_{i}^{\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(t_{i}\cdot v_{j}^{\top}/\tau)},$
where $\cdot$ denotes the dot product used to calculate the similarity between normalized image and text embeddings, and $\tau$ is a learnable temperature parameter. The overall CLIP loss is:
(3) $\mathcal{L}_{CLIP}=\frac{1}{2}(\mathcal{L}_{I\rightarrow T}+\mathcal{L}_{T\rightarrow I}).$
CLIP encodes all semantic content from images and texts into a global representation during its training process, which constrains its capability for compositional understanding. While recent works (Yuksekgonul et al., 2022; Huang et al., 2024a; Zhang et al., 2024a) have sought to enhance this ability by introducing hard negative samples to promote the learning of local semantics, our analysis reveals experimental results demonstrate that such enhancements often degrade the model’s inherent general comprehension capabilities. Consequently, this paper focuses on improving compositional understanding without compromising the original general comprehension abilities of the model.
Table 1. Detailed descriptions of the generated high-quality negative samples.
Negative types Rewrite rule
Intra-sentence reshuffle Subtype1 Reshuffle the overall order of the sentence
Subtype2 Swap the nouns in the sentence
Subtype3 Swap the adjectives in the sentence
Minimal semantic substitution Subtype4 Replace the adjectives in the sentence
Subtype5 Replace the nouns in the sentence
3.2. LLM-Driven Negative Caption Generation
To improve compositional understanding of CLIP, we first leverage the in-context learning capability of Large Language Models (LLMs) to construct about 2M high- quality negative captions across multiple types. Previous methods for generating compositional negative captions can be classified into two categories: rule-based generation (Yuksekgonul et al., 2022; Huang et al., 2024a) and unmasking-based generation (Zhang et al., 2024a; Doveh et al., 2023b, a). However, rule-based approaches are limited in their ability to produce complex compositional negatives, such as those involving minimal semantic substitutions. In contrast, unmasking-based methods can generate diverse negatives but often fail to avoid producing hard positive samples such as “the cat is eating the food” $\rightarrow$ “the cat is eating the hot food”.
To address these limitations, we propose an LLM-driven generation method to enhance and standardize the generation of compositional negative captions. Specifically, we categorize negative captions into two primary types: (1) intra-sentence reshuffling, which improves the VLM’s sensitivity to sentence order, and (2) minimal semantic substitution, which enhances the VLM’s ability to discern subtle semantic variations. For intra-sentence reshuffling, we examine three subtypes: (a) full sentence reordering, (b) noun swapping, and (c) adjective swapping. For minimal semantic substitution, we investigate two subtypes: (a) adjective replacement and (b) noun replacement. Detailed descriptions of each negative sample type are presented in Table 1. As shown in Figure 2, we initially generate high-quality negative samples to fully leverage the in-context learning capabilities of LLMs. For each type, we first use ChatGPT4-Turbo to generate 200 rewritten examples according to rewritten rules, then manually select the 50 highest-quality examples through rigorous evaluation. Building upon these examples, we employ the Llama-3.1-8B-Instruct model [^1] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct to generate large-scale compositional negatives. To ensure semantic divergence from original sentences and avoid hard positives, we incorporate explicit filtering constraints in the prompts (e.g., “generated sentences must exhibit distinct semantics”). Complete prompt templates and representative samples are provided in the appendix.
3.3. DeGLA Framework
Global Alignment. Building upon the base setting (Yuksekgonul et al., 2022), we integrate negative captions into the Equation  1 for global alignment, yielding the augmented image-to-text contrastive:
(4) $\mathcal{L}_{I\rightarrow T}=-\log\frac{\exp(v_{i}\cdot t_{i}^{\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\left(\exp(v_{i}\cdot t_{j}^{\top}/\tau)+\sum_{k=1}^{\mathcal{K}}\exp(v_{i}\cdot\bar{t}_{i,k}^{\top}/\tau)\right)},$
where $\bar{t}_{j,k}$ denotes the $k$-th negative text related to the $j$-th text. In this work, we set $\mathcal{K}=4$ to denote the number of negative samples paired with each positive sample during training, due to the merging of two subtypes, as detailed in the appendix. Then we utilize the hard negative-aware image-text contrastive loss as our base loss which can be formulated as:
(5) $\mathcal{L}_{base}=\frac{1}{2}(\mathcal{L}_{I\rightarrow T}+\mathcal{L}_{T\rightarrow I}),$
However, directly incorporating hard negative text into global contrastive learning significantly compromises the model’s general capabilities, as indicated in Section 4. In this paper, we introduce a self-distillation mechanism within the global alignment framework to mitigate the catastrophic forgetting of pre-trained knowledge. To enhance model robustness and prevent rapid forgetting of pre-trained knowledge during training, we employ an Exponential Moving Average (EMA) strategy to update the weights of the frozen image encoder $\mathcal{E}_{I}^{*}$ and text encoder $\mathcal{E}_{T}^{*}$:
(6) $\mathcal{E}_{I}^{*}$ $=\alpha\mathcal{E}_{I}^{*}+(1-\alpha)\mathcal{E}_{I},$
$\mathcal{E}_{T}^{*}$ $=\alpha\mathcal{E}_{T}^{*}+(1-\alpha)\mathcal{E}_{T},$
where the hyperparameter $\alpha$ controls the update speed of the frozen model’s parameters. Given a batch of image-text pairs $\{I_{i},T_{i},\bar{T}_{i}\}$, we obtain the image and text embeddings $\{v_{i},t_{i},\bar{t}_{i}\}$ from the learnable encoders ($\mathcal{E}_{I},\mathcal{E}_{T}$) and $\{v_{i}^{*},t_{i}^{*},\bar{t}_{i}^{*}\}$ from the frozen EMA encoders ($\mathcal{E}_{I}^{*},\mathcal{E}_{T}^{*}$), respectively. Following L2 normalization, we compute the squared distance between the image-text embeddings from the learnable and frozen EMA encoders to quantify the knowledge discrepancy between the fine-tuned and pre-trained models:
(7) $\mathcal{L}_{Distill}=\sum_{i=1}^{\mathcal{B}}\left(\left\|v_{i}-v_{i}^{*}\right\|_{2}^{2}+\left\|t_{i}-t_{i}^{*}\right\|_{2}^{2}+\sum_{k=1}^{\mathcal{K}}\left\|\bar{t}_{i,k}-\bar{t}_{i,k}^{*}\right\|_{2}^{2}\right).$
Under the constraint of Equation 7, the image-text embeddings undergo subtle refinements within the pre-trained representation space, thereby preserving the model’s general capabilities from excessive degradation.
Figure 3. The proposed training framework of DeGLA.
Table 2. Results (%) on VALSE. The best results are marked in bold, and the second-best results are underlined. The number represents the improvement of our method compared to the CE-CLIP.
Model #Params Existence Plurality Counting Sp.rel. Actions Coreference Foil-it! Avg.
quantifiers number relations repl. actant swap standard clean
BLIP(Li et al., 2022a) 583M 86.3 73.2 68.1 71.5 77.2 61.1 53.8 48.2 93.8 70.0
BEIT3(Wang et al., 2022) 1.9B 77.4 74.6 68.8 74.0 86.7 65.2 50.0 44.2 96.0 70.4
BLIP2(Li et al., 2023b) 3.4B 55.5 71.5 66.0 62.4 83.6 51.6 48.6 51.9 95.9 65.4
MiniGPT-4(Zhu et al., 2023) ¿9B 65.5 72.5 67.4 68.4 83.2 58.8 52.6 51.0 95.8 68.4
Hard Negative based method
XVLM-coco(Zeng et al., 2021a) 216M 83.0 75.6 67.5 70.2 73.8 68.6 46.4 49.6 94.8 69.5
CE-XVLM(Zhang et al., 2024a) 216M 83.5 72.8 72.1 68.7 71.8 69.1 51.0 46.8 93.8 70.8
[…]
CyCLIP(Goel et al., 2022) 151M 69.3 58.3 61.0 66.4 78.1 72.0 53.2 51.6 88.8 65.5
NegCLIP(Yuksekgonul et al., 2022) 151M 76.8 72.0 65.2 72.7 81.6 84.8 58.9 54.8 91.8 71.7
Structure-CLIP (Huang et al., 2024a) 151M 75.6 67.1 62.0 68.2 80.4 88.3 44.5 58.7 91.2 69.1
CE-CLIP(Zhang et al., 2024a) 151M 78.6 77.6 64.3 74.0 81.2 88.5 54.9 52.9 93.7 72.2
DeGLA (ours) 151M 82.4 73.8 68.3 75.3 82.6 88.8 58.5 54.8 93.8 74.1(+1.9)
Table 3. Results(%) on SugarCrepe. Vera and Grammar are text-only models. The best results are marked in bold, and the second-best results are underlined. The number represents the improvement of our method compared to the CE-CLIP.
Model REPLACE SWAP ADD
Object Attribute Relation Avg. Object Attribute Avg. Object Attribute Avg.
Human 100.0 99.0 97.0 98.7 99.0 100.0 99.5 99.0 99.0 99.0
Vera (Liu et al., 2023) 49.4 49.6 49.1 49.4 49.4 49.2 49.3 49.4 49.6 49.5
Grammar (Morris et al., 2020) 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0
BLIP2 (Li et al., 2023b) - - - 86.7 - - 69.8 - - 86.5
Hard Negative based method
CLIP (Radford et al., 2021b) 90.9 80.0 69.2 80.2 62.7 61.4 64.0 77.2 68.2 72.7
NegCLIP (Yuksekgonul et al., 2022) 92.7 85.9 76.5 85.0 75.3 75.2 75.4 88.8 82.8 85.8
Structure-CLIP (Huang et al., 2024a) 91.4 85.0 74.4 83.6 72.7 80.5 76.6 85.5 81.1 83.3
CE-CLIP (Zhang et al., 2024a) 93.1 88.8 79.0 87.0 72.8 77.0 74.9 92.4 93.4 92.9
DeGLA (ours) 94.5 92.6 84.2 90.5(+3.5) 81.6 82.1 81.9(+6.9) 93.8 95.7 94.8(+1.9)
Table 4. Results (%) on ARO. The best results are marked in bold, and the second-best results are underlined. The number represents the improvement of our method compared to the CE-CLIP.
Model Relation Attribute COCO-order Flickr-order Avg.
BLIP(Li et al., 2022a) 59.0 88.0 - - -
BEIT3(Wang et al., 2022) 60.6 74.6 - - -
BLIP2(Li et al., 2023b) 41.2 71.3 - - -
MiniGPT-4(Zhu et al., 2023) 46.9 55.7 - - -
Hard Negative based method
CLIP (Radford et al., 2021b) 59.2 62.9 48.4 59.1 57.4
CyCLIP(Goel et al., 2022) 59.1 65.4 - - -
NegCLIP(Yuksekgonul et al., 2022) 80.4 70.5 86.9 90.5 82.1
Structure-CLIP (Huang et al., 2024a) 81.8 80.5 81.7 83.9 82.0
CE-CLIP(Zhang et al., 2024a) 83.9 76.4 80.9 83.7 81.2
DeGLA (ours) 81.6 74.3 93.8 94.7 86.1(+4.9)
Local Alignment. In addition to the global alignment of $\mathcal{L}_{\text{base}}$, to further enhance compositional understanding, we propose a Local Alignment, which includes Image-Grounded Contrast Loss(IGC) and Text-Grounded Contrast Loss(TGC). We initially introduce the Image-Grounded Contrast loss to attract image embeddings towards positive text embeddings and repel them from negative text embeddings in the feature space (as shown in Figure 3). Specifically, given an image-text pair $(I_{i},T_{i})$ and corresponding hard negative texts $\bar{T}_{i}$, the IGC loss $\mathcal{L}_{IGC}$ is defined as follows:
(8) $\mathcal{L}_{IGC}=-\log\frac{\exp(v_{i}\cdot t_{i}^{\top}/\tau)}{\exp(v_{i}\cdot t_{i}^{\top}/\tau)+\sum_{k=1}^{\mathcal{K}}\exp(v_{i}\cdot\bar{t}_{i,k}^{\top}/\tau)}),$
where $\tau$ is a temperature parameter that controls the sharpness of the distribution. Switching to the textual perspective, hard negative texts, which are minor modifications of the originals, remain proximate to the positive texts within the feature space of the CLIP text encoder. This closeness may lead to mismatches during image embedding alignment with positive text embeddings, potentially impairing compositional understanding. Different from the Image-Grounded Contrast loss, the Text-Grounded Contrast (TGC) loss $\mathcal{L}_{TGC}$ operates solely within the text modality. Its function is to enhance the text encoder’s ability to more effectively discriminate between positive and negative texts, thus improving the compositional understanding of vision language models. The $\mathcal{L}_{TGC}$ is formulated as:
(9) $\mathcal{L}_{TGC}=-\log\frac{\exp(t_{i}\cdot t_{i}^{*\top}/\tau)}{\exp(t_{i}\cdot t_{i}^{*\top}/\tau)+\sum_{k=1}^{\mathcal{K}}\exp(t_{i}\cdot\bar{t}_{i,k}^{\top}/\tau)},$
where $t_{i}^{*}$ represents the progressively frozen text embedding obtained from the frozen EMA text encoder $\mathcal{E}_{T}^{*}$. This embedding is used as an anchor in local alignment for two main reasons: (1) It acts as the positive sample in contrastive learning, enabling the model to discriminate between positive and compositional negative samples; and (2) It mitigates potential overfitting of the text encoder to the feature space of the fine-tuning dataset. Finally, the overall loss function is defined as:
(10) $\mathcal{L}_{all}=\mathcal{L}_{Base}+\lambda_{1}\mathcal{L}_{IGC}+\lambda_{2}\mathcal{L}_{TGC}+\lambda_{3}\mathcal{L}_{Distill}.$
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are loss weights to balance the influence of different loss functions.
4. Experiments
Table 5. Zero-shot classification performance on 11 datasets. The best results are marked in bold, and the second-best results are underlined. The number represents the improvement of our method compared to the CE-CLIP.
Model CIFAR10 CIFAR100 Food101 Pets Flowers SUN397 Cars DTD Caltech101 Aircraft ImageNet Avg.
Pretrained model
CLIP (Radford et al., 2021b) 86.5 61.0 78.5 79.6 58.4 59.9 48.8 38.7 86.3 15.3 57.9 61.0
Hard negative based method
NegCLIP (Yuksekgonul et al., 2022) 86.1 59.9 72.1 78.7 53.9 56.8 43.5 37.7 84.3 11.6 54.0 58.1
Structure-CLIP (Huang et al., 2024a) 76.8 47.4 55.1 61.4 31.3 48.3 16.4 29.4 71.0 7.6 37.3 43.8
CE-CLIP (Zhang et al., 2024a) 80.5 54.1 57.6 59.0 30.1 49.2 22.8 27.6 74.4 9.1 38.1 45.7
DeGLA (ours) 86.5 59.5 75.6 76.0 52.8 59.5 45.7 38.1 84.0 14.1 54.5 58.7(+13.0)
Table 6. Liner probe performance on 11 datasets. The best results are marked in bold, and the second-best results are underlined. The number represents the improvement of our method compared to the CE-CLIP.
Model CIFAR10 CIFAR100 Food101 Pets Flowers SUN397 Cars DTD Caltech101 Aircraft ImageNet Avg.
Pretrained model
CLIP (Radford et al., 2021b) 95.0 80.1 88.5 89.3 94.6 74.1 80.8 73.6 90.5 44.8 74.3 80.5
Hard negative based method
NegCLIP (Yuksekgonul et al., 2022) 94.6 80.0 86.1 89.6 93.9 72.9 78.8 72.9 90.0 43.2 72.9 79.5
Structure-CLIP (Huang et al., 2024a) 91.9 75.5 81.2 86.2 89.6 69.0 67.4 67.7 65.2 37.7 67.7 72.7
CE-CLIP (Zhang et al., 2024a) 94.3 78.5 84.3 88.1 92.6 71.0 74.1 71.8 88.3 39.6 70.7 77.6
DeGLA (ours) 95.1 80.5 86.7 89.5 94.6 74.0 78.8 73.0 89.6 43.5 73.4 79.9(+2.3)
4.1. Implement details
Training Setup. To ensure a fair comparison with previous studies (Yuksekgonul et al., 2022; Zhang et al., 2024a), we employ LLaMA3.1-instruct-8B (Grattafiori et al., 2024) to generate compositional negative samples from the MSCOCO dataset, facilitating direct comparisons with NegCLIP (Yuksekgonul et al., 2022) and CE-CLIP (Zhang et al., 2024a). We utilize the CLIP-ViT/B-32 model as the foundation vision-language model, initializing it with pretrained weights from CLIP (Radford et al., 2021b). The model is fine-tuned on 8 NVIDIA V100 (32G) GPUs for 5 epochs using a batch size of 256, consistent with the protocols of previous works (Zhang et al., 2024a; Yuksekgonul et al., 2022; Doveh et al., 2023b). We employ AdamW as the optimizer, initialized with a learning rate of $1\times 10^{-6}$ and a weight decay of 0.1. The parameters $\alpha$, $\beta_{1}$, and $\beta_{2}$ are set to 0.9996, 0.9, and 0.98, respectively. We perform a hyperparameter search for $\lambda_{1},\lambda_{2},\lambda_{3}$, with optimal values of $\lambda_{1}=0.1$, $\lambda_{2}=0.1$, and $\lambda_{3}=0.005$.
Figure 4. Zero-shot image-text retrieval performance comparison on MSCOCO and Flickr30k.
Evaluation Setup. To comprehensively evaluate DeGLA, we conduct comparative experiments on three compositional reasoning benchmarks: ARO (Yuksekgonul et al., 2022), VALSE (Parcalabescu et al., 2021), and SugarCrepe (Hsieh et al., 2024). To verify that DeGLA retains general capabilities, we further assess its performance on zero-shot classification, liner probe, and retrieval tasks. For zero-shot classification, we evaluate on eleven datasets: CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), Food101 (Bossard et al., 2014), Oxford Pets (Parkhi et al., 2012), Flowers102 (Nilsback and Zisserman, 2008), SUN397 (Xiao et al., 2010), Stanford Cars (Krause et al., 2013), DTD (Cimpoi et al., 2014), Caltech101 (Fei-Fei et al., 2004), FGVC-Aircraft (Maji et al., 2013), and ImageNet (Deng et al., 2009). The datasets used for evaluating the linear probe is the same as that used for zero-shot classification. For zero-shot image-text retrieval, we report results on MSCOCO (Chen et al., 2015) and Flickr30k (Plummer et al., 2015).
We compare DeGLA with three kinds of models: (1)Text-only models, including vera (Liu et al., 2023) and Grammar (Morris et al., 2020). (2) State-of-the-art generative vision-language models, including BLIP (Li et al., 2022a), BLIP-2 (Li et al., 2023b), MiniGPT-4 (Zhu et al., 2023)); (3) High-performance vision-language understanding models, such as BEIT-3 (Wang et al., 2022), XVLM (Zeng et al., 2021b); (4) Specialized compositional improvement methods NegCLIP (Yuksekgonul et al., 2022), CyCLIP (Goel et al., 2022), Structure-CLIP (Huang et al., 2024a), and CE-CLIP (Zhang et al., 2024a).
Figure 5. Performance trade off between compositional reasoning (average performance on VALSE, SugarCrepe, and ARO benchmarks) and zero-shot classification.
(a) Ablation of $\lambda_{1},\lambda_{2}$
(b) Ablation of $\lambda_{3}$
(c) Ablation of hard negative types.
Figure 6. Ablation of the different loss weights and hard negative types.
4.2. Compositional Reasoning
Performance on VALSE. We first evaluate DeGLA on VALSE, a benchmark specifically designed to assess pre-trained vision-language models’ sensitivity to foiled instances. This benchmark systematically tests fundamental linguistic phenomena across visual and linguistic modalities, including existence, plurality, counting, spatial relations, actions, and entity coreference. As indicated in Table 2, DeGLA outperforms other models in existence, counting, and spatial relations, and ranks second in plurality, actions replacement, coreference standard, and foil-it. Compared to CLIP, DeGLA exhibits an 8.8% average performance improvement, demonstrating that our approach more effectively leverages hard negatives through local and global alignment strategies. Thanks to the diverse negative samples generated in this study and the employment of both image-grounded and text-grounded contrastive learning, DeGLA significantly surpasses CE-CLIP, achieving an overall performance increase of 1.9%.
Performance on SugarCrepe. SugarCrepe is a benchmark designed to reduce language bias in existing benchmarks and provide a more accurate assessment of a model’s compositional understanding. As indicated in Table 3, DeGLA achieves new state-of-the-art results across all metrics. DeGLA shows significant improvements over the CLIP model, with increases of 10.3% on Replace, 17.9% on Swap, and 22.1% on Add. These results suggest that DeGLA effectively distinguishes between positive and negative samples through local alignment. Compared to CE-CLIP, our method achieves performance enhancements of 3.5%, 6.9%, and 1.9% on Replace, Swap, and Add, respectively. This substantial improvement is attributed to two primary factors: First, the negative sample generation method we introduced leverages Large Language Models (LLMs) to produce high-quality negative text descriptions, reducing the generation of noisy and challenging positive samples. Second, we implement image-grounded and text-grounded contrast mechanisms that enhance the model’s discriminative capabilities by drawing positive sample pairs closer and pushing negative sample pairs further apart in the feature space, thus improving the model’s compositional understanding.
Performance on ARO. We present the performance of our proposed DeGLA on the ARO benchmark in Table 4. DeGLA achieves substantial improvements, registering an average performance increase of 28.7% over CLIP and 4.9% over CE-CLIP. It is noteworthy that while DeGLA exhibits significant overall performance gains, it still trails CE-CLIP and Structure-CLIP in the domains of relations and attributes. The primary reason for this discrepancy is that our negative sample generation method prioritizes the production of a diverse array of negative samples, rather than focusing specifically on enhancing the understanding of relations and attributes.
Table 7. Ablation of different components. CN: compositional negatives. SD: self-distillation.
Model CN IGC TGC SD ARO ZS-Avg.(11)
CLIP (Radford et al., 2021b) 57.4 61.0
CE-CLIP (Zhang et al., 2024a) 81.2 45.7
DeGLA(ours) 80.8 55.4
84.8 55.1
82.2 57.2
86.2 56.9
86.1 58.7
4.3. General Understanding
Zero-shot classification. In Table 5, we present the zero-shot classification performance across 11 datasets. Notably, while Structure-CLIP and CE-CLIP exhibit robust compositional understanding, they significantly reduce the model’s original general capabilities, with average accuracies decreasing by 17.2% and 15.3%, respectively, compared to CLIP. In contrast, our proposed DeGLA employs a self-distillation constraint module for global alignment, effectively minimizing the loss of general capability. Our method shows a 13.0% improvement in average accuracy over CE-CLIP, while substantially preserving the model’s inherent general capabilities. As illustrated in Figure 5, DeGLA achieves a more effective balance between compositional reasoning and general comprehension capabilities compared to other methods.
Linear probe. In Table 5, we detail the linear probe performance across 11 datasets. Consistent with the zero-shot classification results, Structure-CLIP and CE-CLIP significantly reduce the model’s original general capabilities, with average accuracies decreasing by 7.8% and 2.9%, respectively, compared to CLIP. Conversely, our proposed DeGLA model not only demonstrates superior compositional understanding relative to CE-CLIP but also achieves a 2.3% average performance improvement in linear probe tasks across these datasets.
Figure 7. The case study between CLIP and DeGLA. Green ✔ indicates true caption, while red ✗ indicates false caption. The bar chart represents the model’s prediction results.
Zero-shot image text retrieval. We compare the zero-shot retrieval performance of DeGLA, Structure CLIP, and CE-CLIP on the MSCOCO and Flickr30k datasets. As illustrated in Figure 4, DeGLA surpasses previous models on both datasets, demonstrating superior performance in standard retrieval tasks. This finding is consistent with our earlier experimental results, further substantiating DeGLA’s excellence in both general capabilities and compositional reasoning.
4.4. Analysis
Trade-off Analysis. As illustrated in Figure 5, we analyze the trade-off between general capabilities and compositional reasoning. For the evaluation of compositional reasoning, we utilize the average scores from three benchmarks. Typically, enhancements in compositional reasoning are accompanied by declines in general capabilities, as demonstrated by CE-CLIP, which, despite significant gains in compositional reasoning compared to CLIP, suffers a substantial reduction in general capabilities. In contrast, through its meticulously designed data generation pipeline and training framework, DeGLA achieves an optimal balance. Compared to the baseline CE-CLIP, DeGLA records a 13% improvement in zero-shot classification and an average enhancement of 3.5% in compositional reasoning tasks, further underscoring its superiority.
Ablation of different components. To validate the efficacy of the proposed method, we perform comprehensive ablation studies on key components using the ARO benchmark in Table 7. By combining all modules, DeGLA achieves the most balanced performance, attaining an 86.1% average score on ARO with 4.9% higher than the previous SOTA (CE-CLIP). For zero-shot classification, it outperforms CE-CLIP by 13.0%, demonstrating superior generalization. Analysis of component contributions reveals that integrating high-quality negative samples increases ARO performance by 23.4% but decreases zero-shot classification performance by 5.6%. The local alignment losses, IGC and TGC, improve ARO performance by 4.0% and 1.4%, respectively. After Combining further boosting ARO performance by 5.4%, while zero-shot classification still decreases by 4.1%. Notably, TGC mitigates declines in general capability by using frozen text embeddings from the EMA text encoder, ensuring minimal deviation from the pretrained CLIP model during training. The introduction of a self-distillation mechanism significantly enhances general capability, improving zero-shot classification by 1.8% with a negligible 0.1% loss in ARO performance.
Ablation of Losses. We analyze the impact of the weights $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ of the three losses on the performance of ARO, as illustrated in Figure 6(a) and Figure 6(b). Initially, we identify the optimal combination of ($\lambda_{1},\lambda_{2}$), finding that the model achieves its best performance, 86.1%, with $\lambda_{1}$ = 0.1 and $\lambda_{2}$ = 0.1. Consequently, we adopt these parameters. Subsequent analysis of $\lambda_{3}$ variations reveals that increasing distillation enhances zero-shot classification performance but reduces combinatorial understanding, reflecting the conflicting objectives of self-distillation and compositional reasoning. Ultimately, we select $\lambda_{3}$ = 0.005, as this setting provides a well-balanced model, as indicated by the red dashed circle in the figure.
Ablation of compositional negative texts. To verify the effectiveness of our proposed instruction-based negative generation pipeline, we conduct an ablation analysis on five subtypes of negative texts, as depicted in Figure 6(c). The results show that combining these subtypes achieves optimal performance, thereby validating the pipeline’s effectiveness. This success is attributed to the representativeness of our samples, which effectively enhance the compositional understanding of CLIP.
4.5. Case Study
We provide a case study comparison between CLIP and DeGLA on ARO and SugarCrepe in Figure 7. The results show that the CLIP model exhibits a pronounced “bag-of-words” phenomenon, as indicated in previous work NegCLIP (Yuksekgonul et al., 2022), meaning it struggles with recognizing structural changes in texts, such as swapping the order of targets in a sentence, adding targets, or replacing targets. For instance, CLIP fails to select the correct text from options like “Two ducks swim in a pond with green water.” and “Two swans swim in a pond with green ducks.”. In contrast, DeGLA is capable of correctly distinguishing between them based on the given image. Notably, after training on our diverse compositionality-related hard negative data generated by our proposed pipeline in Section 3.2, DeGLA demonstrates enhanced compositional reasoning understanding compared to pretrained CLIP, such as intra-sentential reshuffle and minimal semantic substitution, thereby underscoring the effectiveness of our data generation pipeline.
5. Conclusion
In this paper, we observe that while previous methods enhance CLIP’s compositional understanding, they often compromise its general capabilities. To overcome this limitation, we introduce a Decoupled Global-Local Alignment (DeGLA) framework, which not only improves compositional understanding but also significantly reduces losses in general capabilities. To optimize the retention of the model’s inherent capabilities, we incorporate a self-distillation mechanism within the global alignment process, aligning the learnable image-text encoder with a frozen teacher model derived from an exponential moving average. For enhancing compositional understanding, we leverage the in-context learning capabilities of Large Language Models (LLMs) to generate approximately 2M high-quality negative captions across five types. We further propose the Image-Grounded Contrast (IGC) loss and Text-Grounded Contrast (TGC) loss to improve vision-language compositionality. Extensive experimental results demonstrate the effectiveness of the DeGLA framework. We hope that our work provides insights into vision-language compositional understanding.
Appendix A Supplementary Material
A.1. Training Details
Details of the hyperparameter configurations are presented in Table 8. An asterisk (*) indicates that the batch size for our positive image-text pairs is 256, and each positive text is paired with four negative texts during training.
Hyperparameters
Batch size 256*
Optimizer AdamW
Weight decay 0.1
Adam $\beta$ (0.9,0.98)
Adam $\epsilon$ 1e-6
Learning rate 1e-6
Learning rate schedule cosine decay
Ema $\alpha$ 0.9996
$(\lambda_{1},\lambda_{2},\lambda_{3})$ (0.1,0.1,0.005)
Epochs 5
Training GPUs 8×8$\times$8 ×V100
Table 8. Detailed hyper-parameters for training DeGLA.
Dataset Classes Train size Test size Evaluation metric
Food101 102 75,750 25,250 accuracy
CIFAR10 10 50,000 10,000 accuracy
CIFAR100 100 50,000 10,000 accuracy
SUN397 397 19,850 19,850 accuracy
Cars 196 8,144 8,041 accuracy
Aircraft 100 6,667 3,333 mean per class
DTD 47 3,760 1,880 accuracy
Pets 37 3,680 3,669 mean per class
Caltech101 101 3,000 5,677 mean-per-class
Flowers 102 2,040 6,149 mean per class
ImageNet 1000 1,281,167 50,000 accuracy
Table 9. List of linear probe datasets with the data distribution and evaluation metrics.
Dataset Test Images Captions per Image Evaluation Protocol
MSCOCO 5,000 5 Image-to-Text & Text-to-Image
Flickr30k 1,000 5 Image-to-Text & Text-to-Image
Table 10. Zero-shot image-text retrieval evaluation settings.
Figure 8. Details of the prompt utilized for generating high-quality examples and negative captions
Benchmark #Examples Composionablity
ARO (Yuksekgonul et al., 2022) 82695 Relation,Attribution,Order
VALSE (Parcalabescu et al., 2021) 8309 Existence, Plurality, Counting, Spatial Relations, Actions,Entity Coreference
SugarCrepe (Hsieh et al., 2024) 5948 Relation,Attribution,Order,Semantic Substitution
Table 11. Details of compositional reasoning benchmarks.
Method Pretrain Finetune Hard negative data Finegrained loss Hard negative type Self distillation
Rule-based Un-masking based LLM-based
CLIP (Radford et al., 2021a)
NegCLIP (Yuksekgonul et al., 2022)
Structure-CLIP (Huang et al., 2024a)
CE-CLIP (Zhang et al., 2024a)
DeGLA (ours)
Table 12. Comparison of different methods.
A.2. Detatils of General Understanding Downstream Datasets
Zero-shot Classification. Following the previous work RWKV-CLIP (Gu et al., 2024), we evaluate the zero-shot classification performance of the models on 11 datasets. The prompt used in zero-shot classification is presented in Table 13.
Linear Probe. The datasets used for the linear probe evaluation are the same as those used for zero-shot classification. Details on each dataset and the corresponding evaluation metrics are provided in Table 9.
Zero-shot Image-Text Retrieval. The detail of zero-shot image-text retrieval benchmark is presented in Table 10. We evaluate retrieval performance on MS-COCO and Flickr30k following standard protocols. For MS-COCO, we use the 1K test set from  (Karpathy and Fei-Fei, 2015), while Flickr30k employs 1,000 images with 5 captions each. In both datasets, we assess bidirectional retrieval (image-to-text and text-to-image) using pre-trained model embeddings without fine-tuning. Performance is measured by Recall@K (R@1, R@5, R@10), representing the percentage of correct matches in the top-K results.
A.3. Details of Compositional Reasoning Benchmarks
In Table 11, we summarize an overview of the three compositional reasoning benchmarks employed in this work.
ARO. ARO (Yuksekgonul et al., 2022) benchmark is used to probe the VLM’s understanding of relations, attributes, and order in image-text data. Specifically, the Relation task involves swapping objects in the text, the Attribute task involves switching attributes in the text, and the Order task involves disrupting the sequence of the entire sentence.
Figure 9. Rewritten examples generated by ChatGPT.
[…]
A figure showing rewritten text examples generated by ChatGPT. Each example consists of an input sentence and its rewritten version.
Figure 10. Hard negative examples generated by LLaMA3.1-8B-instruct.
VALSE. The VASLE (Parcalabescu et al., 2021) benchmark assesses the compositional understanding of Visual Language Models (VLMs) across six dimensions: existence, plurality, counting, spatial relations, actions, and entity coreference. The existence subset employs a single cue to evaluate whether models can discern the presence or absence of specific entities in images. The plurality subset similarly uses a single cue to determine if models can differentiate between singular and plural noun phrases depicted in images, such as “exactly one flower” vs. “some flowers”. The counting subset, which includes balanced, adversarial, and small-number scenarios, tests the models’ ability to confirm the accuracy of a stated number of entities against those shown in the image. The spatial relations subset uses a single cue to assess model performance in identifying different spatial relationships, creating foils by modifying the spatial preposition in the original caption. The actions subset, encompassing action replacement and actant swap settings, evaluates the models’ capability to (i) match the described action with the depicted action, and (ii) accurately identify the agents performing and receiving the action. Finally, the coreference subset examines the models’ proficiency in resolving pronominal references within multimodal contexts, covering both pronouns linked to noun phrases visually grounded in the image and deictic or image-level references.
CIFAR 10 & CIFAR 100
a photo of a {label}. a blurry photo of a {label}. a black and white photo of a {label}. a low contrast photo of a {label}.
a high contrast photo of a {label}. a bad photo of a {label}. a good photo of a {label}. a photo of a small {label}.
a photo of a big {label}. a photo of the {label}. a blurry photo of the {label}. a black and white photo of the {label}.
a low contrast photo of the {label}. a high contrast photo of the {label}. a bad photo of the {label}. a good photo of the {label}.
a photo of the small {label}. a photo of the big {label}.
Food101
a photo of {label}, a type of food.
Caltech101
a photo of a {label}. a painting of a {label}. a plastic {label}. a sculpture of a {label}.
a sketch of a {label}. a tattoo of a {label}. a toy {label}. a rendition of a {label}.
a embroidered {label}. a cartoon {label}. a {label} in a video game. a plushie {label}.
an origami {label}. art of a {label}. graffiti of a {label}. a drawing of a {label}.
a doodle of a {label}. a photo of the {label}. a painting of the {label}. the plastic {label}.
a sculpture of the {label}. a sketch of the {label}. a tattoo of the {label}. the toy {label}.
a rendition of the {label}. the embroidered {label}. the cartoon {label}. the {label} in a video game.
the plushie {label}. the origami {label}. art of the {label}. graffiti of the {label}.
a drawing of the {label}. a doodle of the {label}.
Stanford Cars
a photo of a {label}. a photo of the {label}. a photo of my {label}. i love my {label}!
a photo of my dirty {label}. a photo of my clean {label}. a photo of my new {label}. a photo of my old {label}.
DTD
a photo of a {label} texture. a photo of a {label} pattern. a photo of a {label} thing. a photo of a {label} object.
a photo of the {label} texture. a photo of the {label} pattern. a photo of the {label} thing. a photo of the {label} object.
FGVC Aircraft
a photo of a {label}, a type of aircraft. a photo of the {label}, a type of aircraft.
Flowers102
a photo of a {label}, a type of flower.
Pets
a photo of a {label}, a type of pet.
SUN39
a photo of a {label}. a photo of the {label}.
ImageNet
a bad photo of a {label}. a photo of many {label}. a sculpture of a {label}. a photo of the hard to see {label}.
a low resolution photo of the {label}. a rendering of a {label}. graffiti of a {label}. a bad photo of the {label}.
a cropped photo of the {label}. a tattoo of a {label}. the embroidered {label}. a photo of a hard to see {label}.
a bright photo of a {label}. a photo of a clean {label}. a photo of a dirty {label}. a dark photo of the {label}.
a drawing of a {label}. a photo of my {label}. the plastic {label}. a photo of the cool {label}.
a close-up photo of a {label}. a black and white photo of the {label}. a painting of the {label}. a painting of a {label}.
a pixelated photo of the {label}. a sculpture of the {label}. a bright photo of the {label}. a cropped photo of a {label}.
a plastic {label}. a photo of the dirty {label}. a jpeg corrupted photo of a {label}. a blurry photo of the {label}.
a photo of the {label}. a good photo of the {label}. a rendering of the {label}. a {label} in a video game.
a photo of one {label}. a doodle of a {label}. a close-up photo of the {label}. a photo of a {label}.
the origami {label}. the {label} in a video game. a sketch of a {label}. a doodle of the {label}.
an origami {label}. a low resolution photo of a {label}. the toy {label}. a rendition of the {label}.
a photo of the clean {label}. a photo of a large {label}. a rendition of a {label}. a photo of a nice {label}.
a photo of a weird {label}. a blurry photo of a {label}. a cartoon {label}. art of a {label}.
a sketch of the {label}. a embroidered {label}. a pixelated photo of a {label}. itap of the {label}.
a jpeg corrupted photo of the {label}. a good photo of a {label}. a plushie {label}. a photo of the nice {label}.
a photo of the small {label}. a photo of the weird {label}. the cartoon {label}. art of the {label}.
a drawing of the {label}. a photo of the large {label}. a black and white photo of a {label}. the plushie {label}.
a dark photo of a {label}. itap of a {label}. graffiti of the {label}. a toy {label}.
itap of my {label}. a photo of a cool {label}. a photo of a small {label}. a tattoo of the {label}.
Table 13. Full list of prompts to evaluate the performance of zero-shot classification on 11 visual recognition datasets.
SugarCrepe. The SugarCrepe (Parcalabescu et al., 2021) benchmark encompasses three modalities: Replace, Swap, and Add, enhancing the assessment of visual-linguistic compositional understanding. In the Replace modality, a hard negative is generated by substituting a single atomic concept in a positive caption, which then mismatches with the corresponding image. We classify the replacements according to their conceptual types: REPLACE-OBJ, REPLACE-ATT, and REPLACE-REL. The Swap modality produces hard negatives by interchanging two atomic concepts of the same type within a positive caption, maintaining the original content. This includes SWAP-OBJ and SWAP-ATT categories, excluding relation swaps to avoid generating incoherent text. The Add modality creates hard negatives by introducing a new atomic concept into a positive caption, leading to a discrepancy with the image. It includes ADD-OBJ and ADD-ATT, while relation insertions are excluded due to their implausibility.
A.4. Details of Negative Caption Generation
Following CE-CLIP (Zhang et al., 2024a), we adopt 2014 train split of COCO (Chen et al., 2015) as our base dataset and employ the LLM-driven negative caption generation pipeline to generate compositionally-enhanced data for fine-tuning. The base dataset has 83k images and 414K captions. Through our augmentation process, we generate five hard negative captions per original caption, resulting in a total of 2.07 million hard negative captions.
Prompt Details. In Figure 8, we detail the prompts employed in this study. Notably, the prompts we designed are concise and effective, especially in ensuring that the generated sentences are “hard negative” rather than “hard positive”. To achieve this, we append the instruction “ensuring that the rewritten sentences have significantly different or opposite meanings from the original sentences” at the end of the prompt, which greatly enhanced the quality of both the examples and the final rewritten captions.
Examples Generated by ChatGPT. In Figure 9, we present several representative examples generated by ChatGPT. These examples closely adhere to our rules and serve as high-quality exemplars to guide the large-scale generation of negative samples.
Rewritten Captions. In Figure 10, we present examples of negative samples generated by LLaMA3.1-instruct-8B. These examples demonstrate the high quality of captions produced by the LLM-driven negative caption generation pipeline. The captions adhere accurately to our rewriting rules, ensuring that the generated sentences possess meanings that are distinct from or opposite to the original ones. This efficacy can be attributed to the high-quality examples showcased in Figure 9 and the meticulously designed prompts outlined in Figure 8.